# Sectorial covers over fanifolds

### v2

@@ 5.2. Mirror pair over a fanifold
Up to tailoring $H$, the mirror pair $(H,\partial\mathbf{T}_{\Sigma})$ can be obtained as a special case of mirror pairs $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ over $\Phi$.
Example 5.4 ([GS23, Example 2.11, 4.23]).
Consider the standard embedding
$S^{d}=\{\Sigma^{d+1}_{i=1}x^{2}_{i}=1\}\subset\mathbb{R}^{d+1}$
of the $d$-sphere. Then $\Phi_{\Sigma}=\Sigma\cap S^{d}\subset S^{d}$ carries a canonical fanifold structure. For instance, any vertex $P$ is the intersection of a ray $\rho_{P}\in\Sigma(1)$ and $S^{d}$. To it one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$. Clearly, $\Phi_{\Sigma}$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata.
@@
By Lemma 4.14 we have $\mathbf{T}(\Phi_{\Sigma})=\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})$. As one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$ to each vertex $P\in\Phi_{\Sigma}$, the toric variety $\mathbf{T}(P)$ coincides with $\overline{O(\rho_{P})}$. Let $\sigma_{S^{(k)}}\in\Sigma_{P}$ be the cone corresponding to an exit path in $\Phi_{\Sigma}$ from $P$ to $S^{(k)}$. Then by definition of fanifolds $\mathbf{T}(S^{(k)})$ is the toric variety associated with the fan $\Sigma_{P}/\sigma_{S^{(k)}}\subset M_{\mathbb{R}}/\langle\sigma_{S^{(k)}}\rangle$, which in turn is the toric orbit closure corresponding to the inverse image $\tilde{\sigma}_{S^{(k)}}$ of $\sigma_{S^{(k)}}$ under the quotient $\Sigma\to\Sigma_{P}$. When $S^{(k)}$ runs through the strata of $\Phi_{\Sigma}$, the inverse images $\tilde{\sigma}_{S^{(k)}}$ range over nonzero cones in $\Sigma$. Thus we have
$\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})=\operatorname{colim}_{0\neq\tilde{\sigma}_{S^{(k)}}\in\Sigma}\overline{O(\tilde{\sigma}_{S^{(k)}})}.$
Now, the claim follows from Lemma 5.3. ∎
5.3. Sectorial descent and HMS for very affine hypersufaces
By Lemma 5.5 and Lemma 5.6 the equivalences (5.1) and (4.4) form a commutative diagram
[…]
where the lower equivalence is the global section of the isomorphism $\bar{\pi}_{*}\operatorname{Fuk}_{*}\cong\operatorname{Coh}_{!}\circ\mathbf{T}$ of constructible cosheaves of categories over $\Phi_{\Sigma}$ from Theorem 4.17. Finally, we give an affirmative answer to the conjecture raised in [GS22, Section 2.4].
@@ 5.3. Sectorial descent and HMS for very affine hypersufaces
There is a Weinstein sectorial cover of $H$ lifting the open cover of $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Phi_{\Sigma})$ from Lemma 5.2.
@@
Each piece $\widetilde{\mathbf{W}}(P)$ of our Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a Weinstein sector with skeleton $\mathbb{L}(\Sigma/\rho_{P})$ obtained from $T^{*}\widehat{M}_{P}$ for a vertex $P\in\Phi_{\Sigma}$ by the inductive Weinstein handle attachments. Intersections of $\widetilde{\mathbf{W}}(P)$ are Weinstein sectors-with-sectorial-corners whose skeleta are given by intersections of $\mathbb{L}(\Sigma/\rho_{P})$. Since by Lemma 5.2 the union of $\mathbb{L}(\Sigma/\rho_{P})\cong\mathbb{L}(\Sigma/\rho_{P})\times\partial_{\infty}\rho_{P}$ covers $\partial_{\infty}\mathbb{L}(\Sigma)$ when $\rho_{P}$ runs through the rays in $\Sigma$, the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ gives a lift of the open cover of $\partial_{\infty}\mathbb{L}(\Phi_{\Sigma})$ from Lemma 5.2. ∎
Remark 5.8.
The Weinstein domain $\mathbf{W}(\Phi_{\Sigma})$ may be identified with the Weinstein hypersurface $F\subset\partial_{\infty}D$ which gives the sutured Weinstein manifold $(D,F)$ from [GS22, Corollary 6.2.6]. Due to Lemma 3.11(4), $\widetilde{\mathbf{W}}(\Phi_{\Sigma})$ gives a Weinstein thickening of $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Sigma)=\widetilde{\mathbb{L}}(\Phi_{\Sigma})$ which is the gluing of local Weinstein thickenings of $\mathbb{L}(\Sigma/\rho_{P})$.

### v3

@@ 4.2. Mirror pair over a fanifold
Up to tailoring $H$, the mirror pair $(H,\partial\mathbf{T}_{\Sigma})$ can be obtained as a special case of mirror pairs $(\widetilde{\mathbf{W}}(\Phi),\mathbf{T}(\Phi))$ over $\Phi$. Let $\Phi_{\Sigma}=\Sigma\cap S^{d}\subset S^{d}$ be the fanifold from [GS23, Example 2.11, 4.23] where $S^{d}=\{\Sigma^{d+1}_{i=1}x^{2}_{i}=1\}\subset\mathbb{R}^{d+1}$ is the standard embedding of the $d$-sphere. For instance, any vertex $P$ is the intersection of a ray $\rho_{P}\in\Sigma(1)$ and $S^{d}$. To it one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$. Clearly, $\Phi_{\Sigma}$ is closed and there is no lower dimensional strata which are not adjacent to higher dimensional strata.
@@
By [GS23, Proposition 3.13] we have $\mathbf{T}(\Phi_{\Sigma})=\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})$. As one associates the pair $(\Sigma_{P},M_{P})=(\Sigma/\rho_{P},M/\langle\rho_{P}\rangle)$ to each vertex $P\in\Phi_{\Sigma}$, the toric variety $\mathbf{T}(P)$ coincides with $\overline{O(\rho_{P})}$. Let $\sigma_{S^{(k)}}\in\Sigma_{P}$ be the cone corresponding to an exit path in $\Phi_{\Sigma}$ from $P$ to $S^{(k)}$. Then by definition of fanifolds $\mathbf{T}(S^{(k)})$ is the toric variety for the fan $\Sigma_{P}/\sigma_{S^{(k)}}\subset M_{\mathbb{R}}/\langle\sigma_{S^{(k)}}\rangle$, which in turn is the toric orbit closure corresponding to the inverse image $\tilde{\sigma}_{S^{(k)}}$ of $\sigma_{S^{(k)}}$ under the quotient $\Sigma\to\Sigma_{P}$. When $S^{(k)}$ runs through the strata of $\Phi_{\Sigma}$, the inverse images $\tilde{\sigma}_{S^{(k)}}$ range over nonzero cones in $\Sigma$. Thus from (4.2) we obtain $\operatorname{colim}_{\operatorname{Exit}(\Phi_{\Sigma})^{op}}\mathbf{T}(S^{(k)})=\partial\mathbf{T}_{\Sigma}$. ∎
4.3. Sectorial descent and HMS for very affine hypersurfaces
By Lemma 4.1 and Lemma 4.2 the equivalence (4.1) coincides with (3.3) for $\Phi_{\Sigma}$, i.e. $\operatorname{Fuk}(\widetilde{\mathbf{W}}(\Phi_{\Sigma}))\simeq\operatorname{Coh}(\mathbf{T}(\Phi_{\Sigma}))$. Finally, we give an affirmative answer to the conjecture raised in [GS22, Section 2.5].
@@ 4.3. Sectorial descent and HMS for very affine hypersurfaces
There is a Weinstein sectorial cover of $H$ lifting the open cover of $\operatorname{Core}(H)=\partial_{\infty}\mathbb{L}(\Phi_{\Sigma})$ from [GS22, Corollary 4.8].
@@
Each piece $\widetilde{\mathbf{W}}(P)$ of our Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ is a Weinstein sector with skeleton $\mathbb{L}(\Sigma/\rho_{P})$ obtained from $T^{*}\widehat{M}_{P}$ for a vertex $P\in\Phi_{\Sigma}$ by the inductive Weinstein handle attachments. Intersections of $\widetilde{\mathbf{W}}(P)$ are Weinstein sectors-with-sectorial-corners whose skeleta are given by intersections of $\mathbb{L}(\Sigma/\rho_{P})$. Since by [GS22, Corollary 4.8] the union of $\mathbb{L}(\Sigma/\rho_{P})\cong\mathbb{L}(\Sigma/\rho_{P})\times\partial_{\infty}\rho_{P}$ covers $\partial_{\infty}\mathbb{L}(\Sigma)$ when $\rho_{P}$ runs through the rays in $\Sigma$, the Weinstein sectorial cover $\widetilde{\mathbf{W}}(\Phi_{\Sigma})=\bigcup_{P}\widetilde{\mathbf{W}}(P)$ gives the desired lift. ∎